HIGH RESOLUTION FTIR SPECTROSCOPY OF THE ν_{4} BAND OF BENZENE 1, 3, 5 − d3

\underline{THOMAS A. BLAKE} , ROBERT L. SAMS and STEVEN W. SHARPE, Pacific Northwest National Laboratory, P. O. Box 999, Mail Stop K8-88, 3020 Q Avenue, Richland, WA 99352.

We have recorded the ν_{4} band (Herzberg numbering) of benzene 1, 3, 5 − d3 using high resolution (0.0015 cm^{−1}) Fourier transform infrared spectroscopy. The spectrum was taken with the benzene sample in a temperature stabilized cell (4^{∘}C), 20 cm in length. The interferometer used a mylar beamsplitter and the spectrum was recorded with a liquid helium-cooled silicon bolometer. For the band we have fit the following preliminary spectroscopic constants: ν_{0} = 531.06401(1) cm^{−1}, B′ = 0.171542(2) cm^{−1}, (C′ − B′) − (C′′ − B′′) = 3.48(2) × 10^{−4} cm^{−1}, B′′ = 0.171842(3) cm^{−1}. Band centers of the observed sequences of hot bands will be reported.